# Multiple commutation relations of the quantum affine algebra $U_{q}(\widehat{\mathfrak{gl}}_{N})$, nested Bethe vector and the Gelfand-Tsetlin basis

### v1

@@ 4.1 First specialization
This implies that $\overline{t}^{k}_{k,j-1}=\emptyset$ for $j>k\geq 2$, for all remaining partitions. This effect can be visualized as in Figure 18.
@@ 4.2 Second specialization
$\overline{t}^{1}_{1,1}=\overline{w}_{I_{2}}\cup\cdots\cup\overline{w}_{I_{N}};\qquad\overline{t}^{1}_{1,2}=\dots=\overline{t}^{1}_{1,N-1}=\emptyset.$
@@
$\overline{t}^{j}_{1,k}=\emptyset\qquad\forall j\geq 1,k>1.$
@@
$t^{2}_{22}=\overline{w}_{I_{2}}\cup\cdots\cup\overline{w}_{I_{N}};\qquad t^{2}_{2,3}=\dots=t^{2}_{2,N-1}=\emptyset.$
@@
$\overline{t}^{j}_{jj}=\overline{w}_{I_{j+1}}\cup\cdots\cup\overline{w}_{I_{N}};\qquad\overline{t}^{j}_{k\ell}=\emptyset,\ j\leq k<\ell\leq N.$
@@ 5 Yangian case
is the rational $R$-matrix.
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@ Acknowledgements
K. M. thanks Hitoshi Konno for discussions. This work was partially supported by Grant-in-Aid for Scientific Research (C) 24K06889.
@@
We use the version of the $R$-matrix (2.1), which is obtained by replacing $q$ by $q^{-1}$ in [35, (2.15)].
The quantum minor [35, (2.28)] adopted to our convention is
@@
There is a distinguished commutative subalgebra called the Gelfand-Tsetlin subalgebra generated by $\mathrm{qdet}T^{(j)}(u)$, $j=1,\dots,N$. A basis in which all these elements are simultaneously diagonalized is called the Gelfand–Tsetlin basis. We introduce a partition $J=(J_{1},\dots,J_{N})$ such that
$J_{1}\cup\cdots\cup J_{N}=\{1,\dots,n\},\ \ \ J_{j}\cap J_{k}=\phi\ (j\neq k).$ (A.4)
@@
$\widehat{\xi}_{J}=T_{21}(w_{J_{1}})T_{32}(w_{J_{1}}\cup w_{J_{2}})\cdots T_{NN-1}(w_{J_{1}}\cup\cdots\cup w_{J_{N-1}})e_{N^{n}}.$ (A.5)
@@
Let $\eta$ be a singular vector of weight $\mu(u)=(\mu_{1}(u),\dots,\mu_{k}(u))$ with respect to the subalgebra $U_{q}(\widehat{\mathfrak{gl}}_{n})$. Assume that $\eta$ satisfies $T_{kk}(\alpha)\cdot\eta=0$ for some $\alpha\in\mathbb{C}^{\times}$. Then $T_{k+1,k}(\alpha)\cdot\eta$ is also a singular vector with respect to $U_{q}(\widehat{\mathfrak{gl}}_{n})$, with weight given by
@@
$\widehat{\xi}_{J}^{(j)}=T_{j+1,j}(w_{J_{1}}\cup\cdots\cup w_{J_{j}})\cdots T_{N,N-1}(w_{J_{1}}\cup\cdots\cup w_{J_{N-1}})e_{N^{n}},$ (A.19)
@@
$\widehat{\xi}_{J}=T_{21}(w_{J_{1}})\cdots T_{jj-1}(w_{J_{1}}\cup\cdots\cup w_{J_{j-1}})\widehat{\xi}_{J}^{(j)},$ (A.24)

### v2

@@ 4.2 First specialization
This implies that $\overline{t}^{k}_{k,j-1}=\varnothing$ for $j>k\geq 2$, for all remaining partitions. This effect can be visualized as in Figure 21.
@@ 4.3 Second specialization
$\overline{t}^{1}_{1,1}=\overline{w}_{I_{2}}\cup\cdots\cup\overline{w}_{I_{N}};\qquad\overline{t}^{1}_{1,2}=\dots=\overline{t}^{1}_{1,N-1}=\varnothing.$
@@
$\overline{t}^{j}_{1,k}=\varnothing\qquad\forall j\geq 1,k>1.$
@@
$t^{2}_{22}=\overline{w}_{I_{2}}\cup\cdots\cup\overline{w}_{I_{N}};\qquad t^{2}_{2,3}=\dots=t^{2}_{2,N-1}=\varnothing.$
@@
$\overline{t}^{j}_{jj}=\overline{w}_{I_{j+1}}\cup\cdots\cup\overline{w}_{I_{N}};\qquad\overline{t}^{j}_{k\ell}=\varnothing,\ j\leq k<\ell\leq N.$
@@ 5 Yangian case
is the rational $R$-matrix (Figure 23).
Figure 23: Non-zero matrix elements of the rational $R$-matrix.
@@ Acknowledgements
K. M. thanks Hitoshi Konno for discussions. We thank Jacques H.H. Perk for pointing out relevant references on the nested Bethe ansatz. We are grateful to the referee for the careful reading of our manuscript and for the helpful suggestions. This work was partially supported by Grant-in-Aid for Scientific Research (C) 24K06889.
Declarations
Conflict of Interest Statement
The authors declare that they have no conflicts of interest relevant to the content of this article.
Data Availability Statement
No new data were generated or analyzed in this study; therefore, data sharing is not applicable.
Funding
This research was partially supported by Grant-in-Aid for Scientific Research (C) 24K06889.
@@
We use the version of the $R$-matrix (2.1), which is obtained by replacing $q$ by $q^{-1}$ in [37, (2.15)].
The quantum minor [37, (2.28)] adopted to our convention is
@@
There is a distinguished commutative subalgebra called the Gelfand-Tsetlin subalgebra generated by $\mathrm{qdet}T^{(j)}(u)$, $j=1,\dots,N$. A basis in which all these elements are simultaneously diagonalized is called the Gelfand–Tsetlin basis. We introduce a partition $J=(J_{1},\dots,J_{N})$ as in (1.3).
@@
$\widehat{\xi}_{J}=T_{21}(\overline{w}_{J_{1}})T_{32}(\overline{w}_{J_{1}}\cup\overline{w}_{J_{2}})\cdots T_{NN-1}(\overline{w}_{J_{1}}\cup\cdots\cup\overline{w}_{J_{N-1}})e_{N^{n}}.$ (A.4)
Observe that $e_{N^{n}}$ is obtained in the case $J_{1}=\dots=J_{N-1}=\varnothing$ and $J_{N}=\{1,\dots,n\}$.
@@
Let $\eta$ be a singular vector of weight $\mu(u)=(\mu_{1}(u),\dots,\mu_{k}(u))$ with respect to the subalgebra $U_{q}(\widehat{\mathfrak{gl}}_{n})$. Assume that $\eta$ satisfies $T_{kk}(\alpha)\cdot\eta=0$ for some $\alpha\in\mathbb{C}^{\times}$. Then $T_{k+1,k}(\alpha)\cdot\eta$ is also a singular vector with respect to $U_{q}(\widehat{\mathfrak{gl}}_{n})$, with weight given by, for $u\neq\alpha$,
@@
$\widehat{\xi}_{J}^{(j)}=T_{j+1,j}(\overline{w}_{J_{1}}\cup\cdots\cup\overline{w}_{J_{j}})\cdots T_{N,N-1}(\overline{w}_{J_{1}}\cup\cdots\cup\overline{w}_{J_{N-1}})e_{N^{n}},$ (A.18)
@@
$\widehat{\xi}_{J}=T_{21}(\overline{w}_{J_{1}})\cdots T_{jj-1}(\overline{w}_{J_{1}}\cup\cdots\cup\overline{w}_{J_{j-1}})\widehat{\xi}_{J}^{(j)},$ (A.23)